# Models, Methods and Waveforms for Estimation and Prediction of Doubly Sparse Time-Varying Channels

### v1

@@ III-B Relation to hierarchical sparsity
DS-LTV sparsity is probabilistic while hierarchical sparsity of Definition 2 is deterministic. The two models are nonetheless related: if vectorized to a concatenation of its rows, random matrix $\left[\alpha_{l,q}I_{l,q}\right]_{l,q}$ defines a vector $\boldsymbol{\alpha}\in\mathbb{C}^{(2Q+1)L}$ that consists of $L$ blocks each of size $12Q+$ 1, where on average $s_{\rm d}$ blocks have non-zero entries and where each non-zero block itself is on average $s_{\rm D}$-sparse. To ensure sparsity in a stronger sense i.e., with high probability (as $L,Q,Lp_{\rm d},(2Q+1)p_{\rm D}$ grow), we require that the following assumption hold.
@@
Type-1 and 2 models easily satisfy the CCDF upper bound condition simply by requiring that $\{I_{q}^{(0)}\}_{q}$ in the first and $\{I_{q}^{(l)}\}_{q}$ for any $l$ in the second to be mutually independent (and to thus satisfy $\overline{F}_{S_{\mathrm{D},l}}(m)=\overline{F}_{\mathrm{B}\left(2Q+1,p_{\rm D}\right)}(m),\forall m$). For Type-3 models, $S_{\mathrm{D},l}$ is deterministic and hence its CCDF is trivially upper-bounded. As the following lemma rigorously shows, the mutual independence of $\left\{I_{l}\right\}_{l=0\cdots L-1}$ in Assumption 1 guarantees strong delay domain sparsity while Doppler sparsity is guaranteed in a more explicit manner by the CCDF upper bound.
@@
The proof (which we first gave in [19]) follows from applying Chernoff’s bound to $S_{\rm d}$ evaluated at $s_{\rm d}=(1+\epsilon)Lp_{\rm d}$ and to the $\mathrm{B}\left(2Q+1,p_{\rm D}\right)$ distribution (upper-bounding $S_{\rm D}$ as per Assumption 1) evaluated at $s_{\rm D}=(1+\epsilon)(2Q+1)p_{\rm D}$ with an $\epsilon>0$ chosen as small as needed. ∎
III-C Compressed-sensing estimation of DS-LTV channels using different waveforms
We now turn our attention to the estimation of the DS-LTV channel using practical waveforms and embedded pilots. For that sake, let $\mathbf{s}$ be the $N$-long vector of samples $s_{n}$ at the input of the channel and define $\mathbf{x}$ as the vector of data symbols and embedded channel estimation pilots that modulate the waveform in use to produce the time-domain samples vector $\mathbf{s}$. For all the considered waveforms, we can write the dependence of $\mathbf{s}$ on $\mathbf{x}$ using a modulation matrix $\boldsymbol{\Phi_{\rm tx}}$
@@ III-C Compressed-sensing estimation of DS-LTV channels using different waveforms
For single-carrier modulation (SCM), $\boldsymbol{\Phi}_{\rm tx}=\mathbf{I}_{N}$. For AFDM, $\boldsymbol{\Phi}_{\rm tx}=\boldsymbol{\Lambda}_{c2}\mathbf{F}_{N}\boldsymbol{\Lambda}_{c1}$, with $\boldsymbol{\Lambda}_{c}=\text{diag}(e^{-\imath 2\pi cn^{2}},n=0,\dots,N-1)$ and $\mathbf{F}_{N}$ being the $N$-order discrete Fourier transform (DFT) matrix. For an OFDM grid as the one shown in Figure 3 composed of $N_{\rm s}$ symbols each having $N_{\rm fft}$ sub-carriers and a cyclic prefix of length $N_{\rm cp}$, $\boldsymbol{\Phi}_{\rm tx}=\mathrm{blkdiag}\left(\mathbf{T}_{\rm cp}\mathbf{F}_{N_{\rm fft}}^{\rm H},\ldots,\mathbf{T}_{\rm cp}\mathbf{F}_{N_{\rm fft}}^{\rm H}\right)$ with $\mathbf{T}_{\rm cp}$ being the matrix for cyclic prefix insertion. At the receiver, let $\mathbf{r}$ be the $N$-long vector of the channel output samples $r_{n}$ defined in (4). Depending on the waveform employed, the receiver applies a demodulation matrix $\boldsymbol{\Phi}_{\rm rx}$ to produce the vector of transform domain samples $\mathbf{y}$. For both SCM and AFDM, $\boldsymbol{\Phi}_{\rm rx}=\boldsymbol{\Phi}_{\rm tx}^{\rm H}$. For the OFDM grid of Figure 3, $\boldsymbol{\Phi}_{\rm rx}=\mathrm{blkdiag}\left(\mathbf{F}_{N_{\rm fft}}\mathbf{R}_{\rm cp},\ldots,\mathbf{F}_{N_{\rm fft}}\mathbf{R}_{\rm cp}\right)$ and $\mathbf{R}_{\rm cp}=\left[\mathbf{0},\mathbf{I}_{N_{\rm fft}}\right]$ is the matrix for cyclic prefix removal. Let $\mathcal{P}\subset\llbracket 0\mkern 1.5mu..\mkern 1.5muN-1\rrbracket$ designate the indexes of the received samples associated with $N_{\rm p}$ transform domain pilots, of values $\{{\rm p}_{p}\}_{p=1\cdots N_{\rm p}}$ inserted at indexes $\left\{m_{p}\right\}_{p=1\cdots N_{\rm p}}$ within the vector $\mathbf{x}$ and each surrounded with a waveform-dependent number of zero guard samples. For SCM (Fig. 2),
@@
As for OFDM, let $m_{p_{\mathrm{t}}}\in\llbracket 0\mkern 1.5mu..\mkern 1.5muN_{\rm s}-1\rrbracket$ be the time domain position of the $p$-th pilot and $m_{p_{\mathrm{f}}}\in\llbracket 0\mkern 1.5mu..\mkern 1.5muN_{\rm fft}-1\rrbracket$ be its frequency domain position. Then (Fig. 3)
[…]
While for AFDM, it holds [19] that
@@
The samples related to the $p$-th pilot symbol thus occupy $12N\lvert c_{1}\rvert(L-1)+2Q+$ 1 DAFT domain indexes. More precisely, and in case $c_{1}$ is negative, then (Fig. 4)
@@
Figure 2: SCM frame composed of data samples and $N_{\rm p}$ pilot symbols, each of the latter surrounded by $12L-$ 1 guard samples.
Figure 3: A frame of $N_{\rm s}$ $N_{\rm fft}$-long OFDM symbols, $N_{\rm p,t}$ of which having $N_{\rm p,f}$ pilot subcarriers
Figure 4: AFDM symbol composed of data samples, $N_{\rm p}$ pilot symbols and their guard samples.
For any of the above waveforms, let $\mathbf{y}_{\rm p}\triangleq[y_{k}]_{k\in\mathcal{P}}$ be the vector of received pilot samples. Inserting (8) and (11) into (4) gives the following signal model for recovery of the hierarchically sparse (per Lemma 1) vector $\boldsymbol{\alpha}$
where $[\mathbf{M}]_{l(2Q+1)+Q+q+1}=\boldsymbol{\Phi}_{\rm rx}\boldsymbol{\Delta}_{q}\boldsymbol{\Pi}^{l}\boldsymbol{\Phi}_{\rm tx}\mathbf{x}_{\rm p}$, $\mathbf{x}_{\rm p}$ is a vector of the same length as $\mathbf{x}$ with entries equal to ${\rm p}_{1},\ldots,{\rm p}_{N_{\rm p}}$ at indexes $\left\{m_{p}\right\}_{p=1\cdots N_{\rm p}}$ and to zero elsewhere, and $\mathbf{w}_{\rm p}$ are the corresponding noise samples. Here, $\mathbf{A}_{\cal P}$ is the $|\mathcal{P}|\times N$ matrix that chooses from the transform domain received vector the entries corresponding to $\mathcal{P}$. $\boldsymbol{\Delta}_{q}=\mathrm{diag}(e^{\imath 2\pi qn},n=0\cdots N-1)$, $\boldsymbol{\Pi}$ is the $N$-order permutation matrix, $\boldsymbol{\Phi}$ is the transform matrix specific to the waveform.
Hierarchical hard thresholding pursuit (HiHTP) [18] has been suggested in the literature for solving hierarchically-sparse recovery problems. When applied to Problem (16) it gives Algorithm 1.
Algorithm 1 HiHTP for compressive sensing of DS-LTV channels
1:  Input: $\mathbf{M}_{\rm p}$, $\mathbf{y}_{\rm p}$, maximum number of iterations $k_{\max}$, $s_{\rm d}$, $s_{\rm D}$
2:  $\hat{\boldsymbol{\alpha}}^{\left(0\right)}=0$, $k=0$
@@
4:  $\Omega^{(k+1)}=L_{s_{\rm d},s_{\rm D}}\left(\boldsymbol{\alpha}^{\left(k\right)}+\mathbf{M}_{\rm p}^{\rm H}\left(\mathbf{y}_{\rm p}-\mathbf{M}_{\rm p}\boldsymbol{\alpha}^{\left(k\right)}\right)\right)$
5:  $\boldsymbol{\alpha}^{\left(k+1\right)}={\arg\min}\left\{\left\|\mathbf{y}_{\rm p}-\mathbf{M}_{\rm p}\mathbf{z}\right\|,\sup\left(\mathbf{z}\right)\subset\Omega^{(k+1)}\right\}$
@@
HiHTP employs at each iteration a hierarchically sparse thresholding operator $L_{s_{\rm d},s_{\rm D}}$. To compute $L_{s_{\rm d},s_{\rm D}}(\mathbf{x})$ for a vector $\mathbf{x}\in\mathbb{C}^{L(2Q+1)}$ first a $s_{\rm D}$-sparse approximation is applied to each one of the $L$ blocks of $\mathbf{x}$ by keeping in each of them the largest $s_{\rm D}$ entries while setting the remaining ones to zero. A $s_{\rm d}$-sparse approximation is next applied to the result by identifying the $s_{\rm d}$ blocks with the largest $l_{2}$-norm. To guarantee the convergence of Algorithm 1 and the recovery of $\boldsymbol{\alpha}$, the following technical assumption is needed.
@@
Under Assumption 1 and for sufficiently large $L$, $Q$, sufficiently small $\delta_{\rm t}$, sufficiently small $\delta_{\rm f}$, $N_{\rm p,t}=\Omega(\frac{1}{\delta_{\rm t}^{2}}\log^{2}\frac{1}{\delta_{\rm t}}\log\frac{s_{\rm D}}{\delta_{\rm t}}s_{\rm D}\log(2Q+1))$ and $N_{\rm p,f}=\Omega(\frac{1}{\delta_{\rm f}^{2}}\log^{2}\frac{1}{\delta_{\rm f}}\log\frac{s_{\rm d}}{\delta_{\rm f}}s_{\rm d}\log L)$, then the HiRIP constant $\delta_{s_{\rm d},s_{\rm D}}$ of the SCM measurement matrix $\mathbf{M}_{\rm p}$ associated with $\mathcal{P}\left(N_{\rm p,t}\right)$ as defined by (12) satisfies $\delta_{s_{\rm d},s_{\rm D}}\leq\delta_{\rm t}$ with probability $1-e^{-\Omega\left(\log{L}\log{\frac{s_{\rm D}}{\delta_{\rm t}}}\right)}$ and the HiRIP constant $\delta_{s_{\rm d},s_{\rm D}}$ of the OFDM measurement matrix $\mathbf{M}_{\rm p}$ associated with $\mathcal{P}=\mathcal{P}\left(N_{\rm p,t},N_{\rm p,f}\right)$ as defined by (13) satisfies $\delta_{s_{\rm d},s_{\rm D}}\leq\delta_{\rm t}+\delta_{\rm f}+\delta_{\rm t}\delta_{\rm f}$ with probability $1-e^{-\Omega\left(\min\left\{\log{L}\log{\frac{s_{\rm d}}{\delta_{\rm t}}},\log{(2Q+1)}\log{\frac{s_{\rm D}}{\delta_{\rm f}}}\right\}\right)}$.
@@
Assume $\boldsymbol{\Phi}_{\rm tx}=\boldsymbol{\Lambda}_{c2}\mathbf{F}_{N}\boldsymbol{\Lambda}_{c1}=\boldsymbol{\Phi}_{\rm rx}^{\rm H}$, $\left|c_{1}\right|=\frac{P}{2N}$ and let $P$ be set as the smallest integer satisfying $(L-1)P+2Q+1\geq s_{\rm d}s_{\rm D}$ and $\mathcal{P}=\mathcal{P}\left(N_{\rm p}\right)$ be defined by (15). Then under Assumptions 1 and 2 and for sufficiently large $L$, $Q$, sufficiently small $\delta$, and $N_{\rm p}=\Omega(\frac{1}{\delta^{2}}\log^{2}\frac{1}{\delta}\log\frac{\log(LP)}{\delta}\log(LP)\log\frac{Q}{P})$, the HiRIP constant $\delta_{s_{\rm d},s_{\rm D}}$ of matrix $\mathbf{M}_{\rm p}$ satisfies $\delta_{s_{\rm d},s_{\rm D}}\leq\delta$ with probability $1-e^{-\Omega\left(\log{\left(2\lceil\frac{Q}{P}\rceil+1\right)}\log{\frac{\log(LP)}{\delta}}\right)}$.
@@
When $P\triangleq 2N|c_{1}|$ is set to $12Q+$ 1, AFDM achieves full diversity [15] and the measurements are non-compressive. The setting $P=1$, on the other hand, is the most compressive. By choosing the value of the theorem between these two extremes, each pilot instance gives in its $(L-1)P+2Q+1$-long guard interval a number of measurements close with high probability to the number $s_{\rm d}s_{\rm D}$ of unknowns. The number $N_{\rm p}$ of pilot instances required to estimate the sparsity support has only a logarithmic growth with respect to both delay and Doppler spreads. This property is to be contrasted with the SCM and OFDM HiRIP result showing first-degree polynomial dependence of $N_{\rm p}$ on $s_{\rm d}$ or $s_{\rm D}$ as stated by Theorem 1.
@@
If $\mathbf{M}_{\rm p}$ satisfies the conditions of Theorems 1 or 2, the sequence $\hat{\boldsymbol{\alpha}}^{(k)}$ defined by Algorithm 1 satisfies $\|\hat{\boldsymbol{\alpha}}^{(k)}-\boldsymbol{\alpha}\|\leq\rho^{k}\|\boldsymbol{\alpha}^{(0)}-\boldsymbol{\alpha}\|+\tau\left\|\mathbf{w}_{\rm p}\right\|$ where $\rho<1$ and $\tau$ are constants defined in [18, Theorem 1].

### v2

@@ III-B Relation to hierarchical sparsity [24]
DS-LTV sparsity is probabilistic while hierarchical sparsity of Definition 2 is deterministic. The two models are nonetheless related: if vectorized to a concatenation of its rows, random matrix $\left[\alpha_{l,q}I_{l,q}\right]_{l,q}$ defines a vector $\boldsymbol{\alpha}\in\mathbb{C}^{(2Q+1)L}$ that consists of $L$ blocks each of size $2Q+1$, where on average $s_{\rm d}$ blocks have non-zero entries and where each non-zero block itself is on average $s_{\rm D}$-sparse. To ensure sparsity in a stronger sense, i.e., with high probability (as $L,Q,Lp_{\rm d},(2Q+1)p_{\rm D}$ grow), we require that the following assumption holds.
@@
Type-1 and 2 models satisfy the CCDF upper-bound condition simply by requiring that $\{I_{q}^{(0)}\}_{q}$ in the first and $\{I_{q}^{(l)}\}_{q}$ for any $l$ in the second be mutually independent (and thus satisfy $\overline{F}_{S_{\mathrm{D},l}}(m)=\overline{F}_{\mathrm{B}\left(2Q+1,p_{\rm D}\right)}(m),\forall m$). For Type-3 models, $S_{\mathrm{D},l}$ is deterministic; hence, its CCDF is trivially upper-bounded. As the following lemma rigorously shows, the mutual independence of $\left\{I_{l}\right\}_{l=0\cdots L-1}$ in Assumption 1 guarantees strong delay domain sparsity, while Doppler sparsity is guaranteed in a more explicit manner by the CCDF upper bound.
@@
The proof (which we first provided in [25]) follows from applying Chernoff’s bound to $S_{\rm d}$ evaluated at $s_{\rm d}=(1+\epsilon)Lp_{\rm d}$ and to the $\mathrm{B}\left(2Q+1,p_{\rm D}\right)$ distribution (upper-bounding $S_{\rm D}$ as per Assumption 1) evaluated at $s_{\rm D}=(1+\epsilon)(2Q+1)p_{\rm D}$ with an $\epsilon>0$ chosen as small as needed. ∎
III-C Theoretical analysis of sparse recovery of DS-LTV channels with different waveforms
The above established link to the paradigm of hierarchical sparsity provides us with theoretical tools for an analytical comparison of LTV channel estimation performance when different practical waveforms are used. The comparison is done assuming embedded pilots [^1] As opposed to the joint channel estimation and data detection approach of [26], we only consider pilot-based channel estimation. While data-aided approaches can improve estimation accuracy, we believe that the relative performance comparison results we obtain assuming pure pilot-based channel estimation will continue to hold under the joint approach (though at a higher level of performance for each of the considered waveforms). For that sake, let $\mathbf{s}$ be the $N$-long vector of samples $s_{n}$ at the input of the channel, and define $\mathbf{x}$ as the vector of data symbols and embedded channel estimation pilots that modulate the waveform in use to produce the time-domain samples vector $\mathbf{s}$. For all the considered waveforms, we can write the dependence of $\mathbf{s}$ on $\mathbf{x}$ using a modulation matrix $\boldsymbol{\Phi_{\rm tx}}$
@@ III-C Theoretical analysis of sparse recovery of DS-LTV channels with different waveforms
For single-carrier modulation (SCM), $\boldsymbol{\Phi}_{\rm tx}=\mathbf{T}_{\rm cp}$ while for AFDM, $\boldsymbol{\Phi}_{\rm tx}=\mathbf{T}_{\rm cpp}\boldsymbol{\Lambda}_{c2}\mathbf{F}_{N}^{\rm H}\boldsymbol{\Lambda}_{c1}$, with $\mathbf{T}_{\rm cp}$ and $\mathbf{T}_{\rm cpp}$ being the matrix for (L1)(L-1)-long CP insertion and chirp-periodic prefix (CPP) [25] insertion, respectively, $\boldsymbol{\Lambda}_{c}=\text{diag}(e^{-\imath 2\pi cn^{2}},n=0,\dots,N-1)$ and $\mathbf{F}_{N}$ being the $N$-order discrete Fourier transform (DFT) matrix. For the OFDM grid shown in Figure 3 and composed of $N_{\rm ofdm,symb}$ symbols, each having $N_{\rm fft}$ sub-carriers and a cyclic prefix of length $N_{\rm cp}$, $\boldsymbol{\Phi}_{\rm tx}=\mathrm{blkdiag}\left(\mathbf{T}_{\rm cp}\mathbf{F}_{N_{\rm fft}}^{\rm H},\ldots,\mathbf{T}_{\rm cp}\mathbf{F}_{N_{\rm fft}}^{\rm H}\right)$. For OTFS, $\boldsymbol{\Phi}_{\rm tx}=\mathbf{T}_{\rm cp}\left(\mathbf{F}_{M_{\rm otfs}}^{\rm H}\otimes\mathbf{I}_{N_{\rm otfs}}\right)$ where $M_{\rm otfs}$ and $N_{\rm otfs}$ are two integers satisfying $M_{\rm otfs}N_{\rm otfs}=N$.
At the receiver, let $\mathbf{r}$ be the $N$-long vector of the channel output samples $r_{n}$ defined in (4). Depending on the waveform employed, the receiver applies a demodulation matrix $\boldsymbol{\Phi}_{\rm rx}$ to $\mathbf{r}$ to produce the vector of transform domain samples $\mathbf{y}$. For SCM, $\boldsymbol{\Phi}_{\rm rx}=\mathbf{R}_{\rm cp}$ while for AFDM, $\boldsymbol{\Phi}_{\rm rx}=\boldsymbol{\Lambda}_{c1}^{\rm H}\mathbf{F}_{N}\boldsymbol{\Lambda}_{c2}^{\rm H}\mathbf{R}_{\rm cpp}$ with $\mathbf{R}_{\rm cp}=\mathbf{R}_{\rm cpp}=\left[\mathbf{0},\mathbf{I}_{N}\right]$ being the matrices for (L1)(L-1)-long CP and CPP removal. For the OFDM grid of Figure 3, $\boldsymbol{\Phi}_{\rm rx}=\mathrm{blkdiag}\left(\mathbf{F}_{N_{\rm fft}}\mathbf{R}_{\rm cp},\ldots,\mathbf{F}_{N_{\rm fft}}\mathbf{R}_{\rm cp}\right)$. For OTFS, $\boldsymbol{\Phi}_{\rm rx}=\left(\mathbf{F}_{M_{\rm otfs}}\otimes\mathbf{I}_{N_{\rm otfs}}\right)\mathbf{R}_{\rm cp}$. Let $\{{\rm p}_{p}\}_{p=1\cdots N_{\rm p}}$ be the $N_{\rm p}$ transform domain pilots, inserted at indexes $\left\{m_{p}\right\}_{p=1\cdots N_{\rm p}}$ within the vector $\mathbf{x}$, each surrounded by a waveform-dependent number of zero guard samples. Define $\mathcal{P}\subset\llbracket 0\mkern 1.5mu..\mkern 1.5muN-1\rrbracket$ as the set of indexes of the received samples associated with these pilots. For SCM, each time domain pilot instance needs a number of guard samples that is at least equal to the maximum delay shift (Fig. 2). We thus set
$\mathcal{P}=\bigcup_{p=1}^{N_{\rm p}}\llbracket m_{p}\mkern 1.5mu..\mkern 1.5mum_{p}+L-1\rrbracket.\quad\text{(SCM)}$ (15)
Note that the above pilot scheme with time domain embedded pilots can also work for waveforms other than SCM. This is the case, for instance, of interleave Frequency Division Multiplexing (IFDM) [27], a waveform with a modulation matrix that is particularly relevant for good BER performance with orthogonal approximate message passing (OMAP) and memory approximate message mapping (MMAP) receivers. Therefore, SCM pilot overhead results, that we provide later on, also apply to such waveforms. As for OFDM, let $m_{p_{\mathrm{t}}}\in\llbracket 0\mkern 1.5mu..\mkern 1.5muN_{\rm ofdm,symb}-1\rrbracket$ be the time domain position of the $p$-th pilot and $m_{p_{\mathrm{f}}}\in\llbracket 0\mkern 1.5mu..\mkern 1.5muN_{\rm fft}-1\rrbracket$ be its frequency domain position. Assuming that $N_{\rm fft}$ has been chosen small enough to render negligible the Doppler effect within the individual symbols of the OFDM grid, then no guard samples are needed in the frequency domain, while the multiple CPs play the role of time domain guard intervals (Fig. 3). Then,
@@
While for AFDM with a chirp parameter $c_{1}$, it holds [25] that an $l$ delay shift and a $q$ Doppler shift of a channel path produce a $q-2Nc_{1}l$ DAFT domain shift. More precisely,
@@
The samples related to the $p$-th pilot symbol thus occupy $2N\lvert c_{1}\rvert(L-1)+2Q+1$ DAFT domain indexes. More precisely, each pilot instance in the DAFT domain should be granted $Q$ guard samples preceding it and $2N\lvert c_{1}\rvert(L-1)+Q$ guard samples following it (Fig. 4). We thus set
$\mathcal{P}=\bigcup_{p=1}^{N_{\rm p}}\llbracket m_{p}-Q\mkern 1.5mu..\mkern 1.5mum_{p}+2N\lvert c_{1}\rvert(L-1)+Q\rrbracket.\quad\text{(AFDM)}$ (18)
Finally, for OTFS, the set $\mathcal{P}$ is defined as the vectorized form of the indexes of the cells marked in blue in Figure 5. Note that, since OTFS is a full-diversity waveform for LTV channels, the dimensions of the set $\mathcal{P}$ are adjusted to the channel delay and Doppler spreads; its cardinality cannot thus be reduced below its $L-1+\min(4Q+1,N_{\rm otfs})\min(2L-1,M_{\rm otfs})$ value if pilot-data orthogonality is to be preserved.
Figure 2: SCM frame composed of data samples and $N_{\rm p}$ pilot symbols, each of the latter surrounded by 2L12L-1 guard samples.
Figure 3: A frame of $N_{\rm ofdm,symb}$ $N_{\rm fft}$-long OFDM symbols, $N_{\rm p,t}$ of which having $N_{\rm p,f}$ pilot subcarriers
Figure 4: An AFDM symbol with $N_{\rm p}$ pilot symbols and their guard samples.
Figure 5: An OTFS symbol composed in the Zak domain of data samples (red), a pilot sample (blue) and guard samples (light blue and red)
For any of the above waveforms, let $\mathbf{y}_{\rm p}\triangleq[y_{k}]_{k\in\mathcal{P}}$ be the vector of received pilot samples. Inserting (9) and (14) into (4) gives the following signal model relating the hierarchically-sparse (per Lemma 1) unknown channel vector $\boldsymbol{\alpha}$ to the measurement vector $\mathbf{y}_{\rm p}$ through a measurement matrix $\mathbf{M}_{\rm p}$:
where $\mathbf{M}]_{l(2Q+1)+Q+q+1}=\boldsymbol{\Phi}_{\rm rx}\boldsymbol{\Delta}_{q}\boldsymbol{\Pi}^{l}\boldsymbol{\Phi}_{\rm tx}\mathbf{x}_{\rm p}$, $\mathbf{x}_{\rm p}$ is a vector of the same length as $\mathbf{x}$ with entries equal to ${\rm p}_{1},\ldots,{\rm p}_{N_{\rm p}}$ at indexes $\left\{m_{p}\right\}_{p=1\cdots N_{\rm p}}$ and to zero elsewhere, and $\mathbf{w}_{\rm p}$ are the corresponding noise samples. Here, $\mathbf{A}_{\cal P}$ is the $|\mathcal{P}|\times N$ matrix that selects from the transform domain received vector the entries corresponding to $\mathcal{P}$. $\boldsymbol{\Delta}_{q}=\mathrm{diag}(e^{\imath 2\pi qn},n=0\cdots N-1)$, $\boldsymbol{\Pi}$ is the $N$-order permutation matrix, $\boldsymbol{\Phi}$ is the transform matrix specific to the waveform. Hierarchical hard thresholding pursuit (HiHTP) [24] has been suggested in the literature for solving hierarchically-sparse recovery problems. When applied to Problem (19) it gives Algorithm 1.
Algorithm 1 HiHTP for compressive sensing of DS-LTV channels
1:Input: Measurement matrix $\mathbf{M}_{\rm p}$, received pilot samples $\mathbf{y}_{\rm p}$, max. number of iterations $k_{\max}$, sparsity levels $s_{\rm d}$, $s_{\rm D}$
@@
[…]
@@
HiHTP employs a hierarchically sparse thresholding operator $L_{s_{\rm d},s_{\rm D}}$ at each iteration. To compute $L_{s_{\rm d},s_{\rm D}}(\mathbf{x})$ for a vector $\mathbf{x}\in\mathbb{C}^{L(2Q+1)}$ a $s_{\rm D}$-sparse approximation is first applied to each of the $L$ blocks of $\mathbf{x}$ by keeping the largest $s_{\rm D}$ entries in each block while setting the remaining ones to zero. A $s_{\rm d}$-sparse approximation is then applied to the result by identifying the $s_{\rm d}$ blocks with the largest $l_{2}$-norm. HiHTP is known to converge, with a geometric convergence rate, provided that the associated HiRIP is small enough [24, Theorem 1]. This applies to Algorithm 1 (as established by Corollary 1 below) provided that the following technical assumption holds.
@@
Under the above assumption, the following two theorems characterize the pilot overhead needed for each of the studied waveforms in order to make the associated HiRIP small enough and hence to guarantee the convergence of Algorithm 1 and the consequent recovery of $\boldsymbol{\alpha}$.
@@
Under Assumption 1 and for sufficiently large $L$, $Q$, sufficiently small $\delta_{\rm t}$, sufficiently small $\delta_{\rm f}$, and
$N_{\rm p,t}=\Omega(\frac{1}{\delta_{\rm t}^{2}}\log^{2}\frac{1}{\delta_{\rm t}}\log\frac{s_{\rm D}}{\delta_{\rm t}}s_{\rm D}\log(2Q+1))\>,$ (20)
$N_{\rm p,f}=\Omega(\frac{1}{\delta_{\rm f}^{2}}\log^{2}\frac{1}{\delta_{\rm f}}\log\frac{s_{\rm d}}{\delta_{\rm f}}s_{\rm d}\log L)\>,$ (21)
then the HiRIP constant $\delta_{s_{\rm d},s_{\rm D}}$ of the SCM measurement matrix $\mathbf{M}_{\rm p}$ associated with $\mathcal{P}\left(N_{\rm p,t}\right)$ as defined by (15) satisfies $\delta_{s_{\rm d},s_{\rm D}}\leq\delta_{\rm t}$ with probability $1-e^{-\Omega\left(\log{L}\log{\frac{s_{\rm D}}{\delta_{\rm t}}}\right)}$ and the HiRIP constant $\delta_{s_{\rm d},s_{\rm D}}$ of the OFDM measurement matrix $\mathbf{M}_{\rm p}$ associated with $\mathcal{P}=\mathcal{P}\left(N_{\rm p,t},N_{\rm p,f}\right)$ as defined by (16) satisfies $\delta_{s_{\rm d},s_{\rm D}}\leq\delta_{\rm t}+\delta_{\rm f}+\delta_{\rm t}\delta_{\rm f}$ with probability $1-e^{-\Omega\left(\min\left\{\log{L}\log{\frac{s_{\rm d}}{\delta_{\rm t}}},\log{(2Q+1)}\log{\frac{s_{\rm D}}{\delta_{\rm f}}}\right\}\right)}$.
@@
Assume $\boldsymbol{\Phi}_{\rm tx}=\boldsymbol{\Lambda}_{c2}\mathbf{F}_{N}\boldsymbol{\Lambda}_{c1}=\boldsymbol{\Phi}_{\rm rx}^{\rm H}$, $\left|c_{1}\right|=\frac{P_{\rm afdm}}{2N}$ and let $P_{\rm afdm}$ be the smallest integer satisfying $(L-1)P_{\rm afdm}+2Q+1\geq s_{\rm d}s_{\rm D}$ and $\mathcal{P}=\mathcal{P}\left(N_{\rm p}\right)$ be defined by (18). Then, under Assumptions 1 and 2 and for sufficiently large $L$, $Q$, sufficiently small $\delta$, and
[…]
the HiRIP constant $\delta_{s_{\rm d},s_{\rm D}}$ of matrix $\mathbf{M}_{\rm p}$ satisfies $\delta_{s_{\rm d},s_{\rm D}}\leq\delta$ with probability $1-e^{-\Omega\left(\log{\left(2\lceil\frac{Q}{P_{\rm afdm}}\rceil+1\right)}\log{\frac{\log(LP_{\rm afdm})}{\delta}}\right)}$.
@@
When $P_{\rm afdm}\triangleq 2N|c_{1}|$ is set to $2Q+1$, AFDM achieves full diversity [20] and the measurements are non-compressive. The setting $P_{\rm afdm}=1$, on the other hand, is the most compressive. By choosing the value of $P_{\rm afdm}$ between these two extremes, as in the theorem, each pilot instance provides in its $(L-1)P_{\rm afdm}+2Q+1$-long guard interval a number of measurements that is close, with high probability, to the number $s_{\rm d}s_{\rm D}$ of unknowns. Furthermore, the theorem states that $N_{\rm p}$, the number of AFDM pilot instances required to estimate the sparsity support has only logarithmic growth with respect to both delay and Doppler spreads. This property is to be contrasted with the first-degree polynomial dependence of $N_{\rm p}$ on $s_{\rm d}$ or $s_{\rm D}$ in the case of SCM and OFDM.
@@
If $\mathbf{M}_{\rm p}$ satisfies the conditions of Theorems 1 or 2, the sequence $\hat{\boldsymbol{\alpha}}^{(k)}$ defined by Algorithm 1 satisfies $\|\hat{\boldsymbol{\alpha}}^{(k)}-\boldsymbol{\alpha}\|\leq\rho^{k}\|\boldsymbol{\alpha}^{(0)}-\boldsymbol{\alpha}\|+\tau\left\|\mathbf{w}_{\rm p}\right\|$ where $\rho<1$ and $\tau$ are constants defined in [24, Theorem 1].